# Quantitative estimates: How well does the discrete Fourier transform approximate the Fourier transform on $\mathbb{R}$?

### v1

@@ 1 Introduction
Let $r>0$ and $0<\alpha\leq 1$. If $f$ and $\hat{f}$ are continuous and satisfy $|f(x)|\lesssim e^{-r|x|^{\alpha}}$ and $|\hat{f}(\xi)|\lesssim(1+|\xi|)^{-b}$ with $b>1$, then, for all $r^{\prime}<r$ and $\beta<b-\frac{1}{2}$,
$E^{[n]}_{h}(f)\lesssim e^{-r^{\prime}(\frac{p}{2})^{\alpha}}\sup_{x\in\mathbb{R}}|f(x)|e^{r|x|}+h^{\beta}\sup_{\xi\in\mathbb{R}}|\hat{f}(\xi)|(1+|\xi|)^{b}\,.$ (6)
For all practical purposes, the length of the sampled interval $p$ is sufficiently large, so that $e^{-r^{\prime}(\frac{p}{2})^{\alpha}}$ is negligible (and may for instance even be close to machine precision). Then the error is dominated by the polynomial term $h^{\beta}$.
In the main theoretical part in Section 3, we will derive more general results by quantifying the decay of a function by means of its membership in a Wiener amalgam space. This is the canonical function class when sampling is involved [9, 10, 12].
Software to plot the discrete Fourier transform commonly performs interpolation of the computed data vector $\operatorname{\mathcal{F}}\!f_{h,n}$. In order to simulate this process, we use three representative interpolation procedures and derive bounds on their deviation from $\hat{f}$ on the real line. One of them is based on the cardinal sine function $\operatorname*{sinc}(x)=\frac{\sin(\pi x)}{\pi x}$ that is widely used in digital signal processing. The computation of $\operatorname{\mathcal{F}}\!f_{h,n}$ gives rise to the approximation of $\hat{f}$ on the real line by
@@
So the error is then $\|\hat{f}-h\sum_{k,j\in[n]}f(hj)e^{-2\pi\mathrm{i}\frac{kj}{n}}\operatorname*{sinc}(p\cdot-k)\|_{L^{2}(\mathbb{R})}$. We will show in Section 5, Theorem 4.1, that this error obeys the same bounds as in Theorems 1.1, 1.2, and 1.3. The proof uses the estimates for $E^{[n]}_{h}(f)$.
@@
Outline: Section 2 is dedicated to numerical experiments using the FFT and fully supports the theoretical results of Theorems 1.1 and 1.3. The full versions of the theoretical results on the approximation rates for spaces of polynomial and sub-exponential decay in time and in frequency are derived in Section 3. The approximation of $\hat{f}$ on the real line is discussed in Section 4. We investigate minimal requirements on $f$ and $\hat{f}$ for convergence in Section 5.
2 Numerical experiments for polynomial decay
Since our priority is the practical impact of the error estimates, we first offer numerical evidence that the error bounds of Theorem 1.1 are sharp for polynomial weights.
@@ 2.1 Polynomial decay in time and in frequency
We use linear combinations of shifts of the cardinal B-splines defined by $B_{1}:={\bf 1}_{[-\frac{1}{2},\frac{1}{2})}$, and
@@
Each $B_{b}$ is a piecewise polynomial function of degree $b-1$ that is $b-2$-times continuously differentiable with support $[-\frac{b}{2},\frac{b}{2}]$, see [20] and also [19, Section 9.1].
@@
so that $c_{2k}=0$. We study the family of functions
$f^{a,b}(x):=\sum_{k\in\mathbb{Z}}c^{a}_{k}\,B_{b}(x-k),\qquad a,b=1,2,\ldots\,,$ (7)
which consists of locally finite sums. Therefore the point evaluations can be evaluated accurately in numerical experiments, and they satisfy $|f^{a,b}(x)|\lesssim(1+|x|)^{-a}$.
To compute the Fourier transform of $f^{a,b}$, we define the Fourier series $u_{a}(\xi):=\sum_{k\in\mathbb{Z}}c^{a}_{k}e^{-2\pi\mathrm{i}k\xi}$. As $\widehat{B_{b}}=\operatorname*{sinc}^{b}$, we observe
@@
The estimate $|\operatorname*{sinc}^{b}(\xi)|\lesssim(1+|\xi|)^{-b}$ implies $|\widehat{f^{a,b}}(\xi)|\lesssim(1+|\xi|)^{-b}$, thus $f^{a,b}$ satisfies the assumptions of Theorem 1.1.
@@
For $a=2,3$ we have $u_{2}(\xi)=|\xi|$ $u_{a}$ for $\xi\in[-\frac{1}{2},\frac{1}{2})$, and $u_{3}(\xi)=\operatorname*{sign}(\xi)\xi^{2}-\frac{1}{2}\xi+\frac{1}{8}$. Thus, we can evaluate the point evaluations of $f^{a,b}$ and $\widehat{f^{a,b}}$ exactly, both analytically and numerically, and the decay is labeled by $a$ and $b$. For the numerical simulations we compute the samples $\widehat{f^{a,b}}(k/p)$ and compare them with the discrete Fourier transform of a sampled version of $f^{a,b}$.
Since the error estimates of Theorem 1.1 work for every parameter $\alpha<a-1/2$ and $\beta<b-1/2$, we will plot the error curves with the values $\alpha=a-\frac{1}{2}$ and $\beta=b-\frac{1}{2}$ and compare our numerical observations with the error rate $n^{-\frac{\alpha\beta}{\alpha+\beta}}$ for $h=n^{-\frac{\alpha}{\alpha+\beta}}$. The infinitesimal difference between $\alpha<a-1/2$ and $\alpha=a-1/2$ is not visible.
To compute the discrete Fourier transform, we choose $n=2^{l}$, $l=10,\ldots,20$ in Theorem 1.1 and then apply the FFT to the samples $f^{a,b}_{h,n}$. The FFT is taken from the Julia package FFTW.jl with $64$-bit floating point arithmetics. The package provides a binding to the C library FFTW [13].
The approximation rates, predicted by Theorem 1.1, are $n^{-\frac{\alpha\beta}{\alpha+\beta}}\asymp h^{\beta}$, see Table 1 for some combinations of $a$ and $b$. To stick to $L^{1}(\mathbb{R})$ functions, we skip $a=1$.
@@
Our numerical experiments as illustrated in Figure 1 align with the theoretical bounds in Theorem 1.1 and suggest that they are sharp.
Figure 1: Logarithmic plots of $E^{[n]}_{h}(f^{a,b})$ for various $a$ and $b$. Points are computed numerically for $n=2^{l}$, $l=10,\ldots,20$. Reference lines represent the theoretical slopes from Theorem 1.1 listed in Table 1.
@@
The closed form expressions for $u_{a}(\xi)=\sum_{k\in\mathbb{Z}}c^{a}_{k}e^{-2\pi\mathrm{i}k\xi}$ are key to our numerical results: Without a closed form expression of $u_{a}$ the series in (7) would need to be truncated for evaluation. The truncation error is accompanied by competing floating point errors due to the summation. We found it difficult to identify a sufficiently accurate truncation that yields reliable numerical estimates for $E^{[n]}_{h}(f^{a,b})$.
2.2 Polynomial decay in frequency
To test Theorem 1.3, we consider $f(x)=e^{-2\pi|x|}$. Its Fourier transform is $\hat{f}(\xi)=\frac{1}{\pi(1+\xi^{2})}$, so that $r=2\pi$, $\alpha=1$, and $|\widehat{f}(\xi)|\lesssim(1+|\xi|)^{-b}$ with $b=2$.
For increasing values of $n$ in numerical experiments with the FFT, we choose successively $p=n^{1/2}$, $p=6n^{1/10}$, and $p=10$, so that $h=n^{-1/2}$, $h=6n^{-9/10}$, and $h=10n^{-1}$, respectively. Note that $p$ is sufficiently large, so that the error bound in Theorem 1.3 is expected to be governed by the polynomial term for a large range of $n$. Hence, we expect to observe the decay rate $h^{\frac{3}{2}}$, which is $n^{-\frac{3}{4}}$, $n^{-\frac{27}{10}}$, and $n^{-\frac{3}{2}}$, respectively. Figure 2 shows the results of our numerical experiments. Again, they match the theoretical prediction perfectly.

### v2

@@ 2.4 Mixed conditions
Let $r>0$ and $0<\alpha\leq 1$. If $f$ and $\hat{f}$ are continuous and satisfy $\sup_{x\in\mathbb{R}}|f(x)|e^{r|x|}<\infty$ and $\sup_{\xi\in\mathbb{R}}|\hat{f}(\xi)|(1+|\xi|)^{b}<\infty$ with $b>1$, then, for all $r^{\prime}<r$ and $\beta<b-\frac{1}{2}$, there is a constant $c>0$ that is independent of $h$, $n$, $p$ such that
(10) $E^{[n]}_{h}(f)\leq c\,(e^{-r^{\prime}p/2}+h^{\beta})\,.$
For all practical purposes, the length of the sampled interval $p$ is sufficiently large, so that $e^{-r^{\prime}p/2}$ is negligible (and usually close to machine precision). Then the error is dominated by the polynomial term $h^{\beta}$.
The above estimate implies a pointwise error of $h^{\beta}$ for $\beta<b-1/2$. The error estimate (4) from [5] induces a bound for the $\ell^{2}$-error $E^{[n]}_{h}(f)$ on the order of $h^{\beta}$ for $\beta=b-\frac{1}{2}$, however, under additional, stringent boundary conditions.
For comparison, we formulate the error estimate for the known case of functions with compact support.
Corollary 2.5.
Assume that $f$ has compact support in $\frac{p}{2},\frac{p}{2}]$ with $p=hn$ and is $m$-times differentiable, then for every $\epsilon>0$ there is a constant $c=c(\epsilon,p)$, such that
$E^{[n]}_{h}(f)\leq c\,n^{-(m-1/2-\epsilon)}\,.$
We note that the pointwise estimate $n^{-m}$ from [11] translates into the estimate $E^{[n]}_{h}(f)\lesssim n^{-(m-1/2)}$. Thus except for the additional $\epsilon$ in the exponent, we recover the well-known results for compactly supported functions. The presence of $\epsilon$ is plausible, because the assumption of differentiability is slightly stronger than the polynomial decay of $\hat{f}$, and we derive Corollary 2.5 from that weaker assumption (and thus obtain a weaker conclusion).
2.5 Interpolation
Software to plot the discrete Fourier transform commonly performs interpolation of the computed data vector $\operatorname{\mathcal{F}}\!f_{h,n}$. The implicit question is then what the samples of $\hat{f}(k/p)$ or their approximations say about the global Fourier transform $\hat{f}$. In order to simulate this process, we use three representative interpolation procedures and derive bounds on their deviation from $\hat{f}$ on the real line. One of them is based on the cardinal sine function $\operatorname*{sinc}(x)=\frac{\sin(\pi x)}{\pi x}$ that is widely used in time-frequency analysis, digital signal processing, and numerical analysis [6, 24, 27, 28]. The computation of $\operatorname{\mathcal{F}}\!f_{h,n}$ gives rise to the approximation of $\hat{f}$ on the real line by
@@ 2.5 Interpolation
We will consider the $L^{2}$ norm, this is, we estimate
$\|\hat{f}-h\sum_{k,j\in[n]}f(hj)e^{-2\pi\mathrm{i}\frac{kj}{n}}\operatorname*{sinc}(p\cdot-k)\|_{L^{2}(\mathbb{R})}\,$,
@@
and will show in Section 5, Theorem 5.1, that this error obeys exactly the same bounds as $E_{h}^{[n]}(f)$ in Theorems 2.1, 2.3, and 2.4.
2.6 Minimal requirements for convergence
Whereas in most applications and in scientific computing it is important to understand how well the FFT approximates the Fourier transform and how fast the convergence happens, it is a fundamental question whether and when the approximations converge at all. Mathematically, we search for minimal conditions on a function. To the best of our knowledge, this problem has not been addressed in the literature.
For the initial formulation, we use the pointwise error and weak polynomial decay conditions.
Proposition 2.6.
If $f$ and $\hat{f}$ satisfy the mild decay conditions $\sup_{x\in\mathbb{R}}|f(x)|(1+|x|)^{1+\epsilon}<\infty$ and $\sup_{\xi\in\mathbb{R}}|\hat{f}(\xi)|(1+|\xi|)^{1+\epsilon}<\infty$ for some $\epsilon>0$, then
(11) $\lim_{p\to\infty}\lim_{h\to 0}\sup_{k\in[n]}\Big|\hat{f}(\tfrac{k}{p})-h\sum_{j\in[n]}f(hj)e^{-2\pi\mathrm{i}\frac{kj}{n}}\Big|=0\,.$
Clearly, these decay conditions are the weakest possible. If $\epsilon=0$, then $f$ need not be integrable and its Fourier transform may not even be defined pointwise.
In Section 6 we will formulate a more general, but technical condition that is related to the validity of the Poisson summation formula, and we will also treat the convergence of the interpolation.
2.7 Other error measures and noisy samples
We will be exclusively interested in the $\ell^{2}$-approximation error $E^{[n]}_{h}(f)$, that is the error made by replacing the Fourier transform by the FFT. Other error norms or the influence of other effects, such as rounding or distortions, can be derived from the main estimates.
(i) Pointwise error estimates follow easily because
(12) $\max_{k\in[n]}\Big|\hat{f}\Big(\tfrac{k}{p}\Big)-h\sum_{j\in[n]}f(hj)e^{-2\pi i\frac{kj}{n}}\Big|\leq\sqrt{p}E^{[n]}_{h}(f)\,$,
and other $\ell^{q}$-norms can be handled similarly. As an example, under the decay assumptions of Theorem 2.1, (8) implies the pointwise error estimate
$\max_{\in[n]}\Big|\hat{f}(\tfrac{k}{p})-h\sum_{j\in[n]}f(hj)e^{-2\pi\mathrm{i}\frac{kj}{n}}\Big|\lesssim n^{-\frac{\beta}{\alpha+\beta}(\alpha-\frac{1}{2})}\,$,
where the last inequality follows for the choice $p^{-\alpha}\asymp n^{-\frac{\alpha\beta}{\alpha+\beta}}$. We do not claim that the bounds for the pointwise error are optimal. Our tools are designed for $E^{[n]}_{h}(f)$ and rely heavily on its time-frequency symmetry, which is not shared by the pointwise error.
(ii) Noise, contaminated function evaluation, or round-off errors are important sources of errors in scientific computing. In each case, instead of the correct samples $f(jh)$ the samples of a distorted function $\widetilde{f}$ or noisy samples $\widetilde{f}(hj)=f(hj)+\epsilon_{j}$ are used as input of the FFT. However, since the approximation (2) is linear and the FFT is unitary, the effect of these errors is benign and remains what it was. Formally, using $\widetilde{f}$ and its samples $\widetilde{f}_{h,n}$ instead of $f_{h,n}$ as input of the FFT, the resulting error is
(13) $\big\|\hat{f}_{\frac{1}{p},n}-\operatorname{\mathcal{F}}\!\widetilde{f}_{h,n}\big\|\leq\big\|\hat{f}_{\frac{1}{p},n}-\operatorname{\mathcal{F}}\!f_{h,n}\big\|+\big\|\operatorname{\mathcal{F}}\!f_{h,n}-\operatorname{\mathcal{F}}\!\widetilde{f}_{h,n}\big\|=E^{[n]}_{h}(f)+\big\|(\widetilde{f}-f)_{h,n}\big\|\,.$
Thus the error splits into the approximation error $E^{[n]}_{h}(f)$, which we are concerned with, and the norm of the initial noise. Consequently, all error bounds remain valid for the distorted function values after adding the noise term $\big\|(\widetilde{f}-f)_{h,n}\big\|$ to the original bound.
2.8 Methods
As in [5] we will use the Poisson summation formula to decompose the approximation error into a time and a frequency component. This allows us to estimate each component separately. In order to make this approach precise, we need to sample a function on a grid and to apply a pointwise version of Poisson’s summation formula. The novelty of our approach is (i) the systematic exploitation of the time-frequency symmetry of the problem, and (ii) the use of so-called Wiener amalgam spaces to describe sampling and decay properties. This class is the canonical function class when sampling is involved [12, 13, 15] and is widely used in Fourier analysis. We believe that it is also very useful and convenient in numerical analysis.
Outline
@@ 2.8 Methods
Section 3 is dedicated to numerical simulations to illustrate that the theoretical results of Theorems 2.1 and 2.4 are best possible. The full versions of the theoretical results on the approximation rates for spaces of polynomial and sub-exponential decay in time and in frequency are derived in Section 4. The approximation of $\hat{f}$ on the real line via interpolation is discussed in Section 5. In Section 6 we investigate under which minimal conditions the error tends to zero.
Acknowledgements
The authors thank Norbert Kaiblinger for valuable suggestions that helped improving the manuscript. We would like to express our gratitude to the referees for their meticulous reading and detailed comments on the manuscript. Their suggestions have helped us immensely to improve the manuscript. A. K. is supported by the FWF project DISCO (PAT4780023).
3 Numerical simulations
We back up the theoretical results with numerical evidence that the polynomial error bounds in Theorem 2.1 are sharp, up to some infinitesimal margin.
It is important to emphasize that the following numerical examples are not intended to represent typical functions for a specific application. The sole purpose of this section is to illustrate that the assumptions and error rates associated with polynomial decay cannot, in general, be improved. Strictly speaking, our theoretical results allow for arbitrary $\alpha<a-\frac{1}{2}$, while the numerics suggest that the rate $\alpha=a-\frac{1}{2}$ still holds. This is what we mean by ‘sharp up to some infinitesimal margin’.
This numerical verification of optimality is nontrivial: it requires functions $f$ for which both $f$ and its Fourier transform $\hat{f}$ can be evaluated with high accuracy, and for which the decay rates in both time and frequency are precisely known. For functions with a standard closed-form expression, at least one of these properties typically fails. We therefore construct suitable functions through a more refined approach below.
@@ 3.1 Polynomial decay in time and in frequency
We use infinite linear combinations of shifts of the cardinal B-splines defined by $B_{1}:={\bf 1}_{[-\frac{1}{2},\frac{1}{2})}$, and
@@
Each $B_{b}$ is a piecewise polynomial function of degree b1b-1 that is b2b-2-times continuously differentiable with support $\frac{b}{2},\frac{b}{2}]$, and its Fourier transform is $\widehat{B_{b}}(\xi)=\big(\tfrac{\sin\pi\xi}{\pi\xi}\big)^{b}=\operatorname*{sinc}^{b}(\xi)$, see [26] and also [24, Section 9.1].
@@
so that $c_{2k}=0$. For integer parameters $a,b\in\mathbb{N}$ we study the family of functions
(14) $f^{a,b}(x):=\sum_{k\in\mathbb{Z}}c^{a}_{k}\,B_{b}(x-k),\qquad a,b=1,2,\ldots\,.$
Every $f^{a,b}$ is a locally finite sum, therefore its point evaluations can be computed accurately in numerical experiments, and they satisfy $\sup_{x\in\mathbb{R}}|f^{a,b}(x)|(1+|x|)^{a}<\infty$.
To compute the Fourier transform of $f^{a,b}$, we define the Fourier series $u_{a}(\xi):=\sum_{k\in\mathbb{Z}}c^{a}_{k}e^{-2\pi\mathrm{i}k\xi}$ and obtain
@@
The estimate $|\operatorname*{sinc}^{b}(\xi)|\lesssim(1+|\xi|)^{-b}$ implies $\sup_{\xi\in\mathbb{R}}|\widehat{f^{a,b}}(\xi)|(1+|\xi|)^{b}<\infty$, thus $f^{a,b}$ satisfies the assumptions of Theorem 2.1.
@@
For $a=2,3,4,5$, tedious calculations lead to periodic functions with period $1$ whose values on [1/2,1/2)[-1/2,1/2) are given by
$u_{2}(\xi)$ $=|\xi|,$
$u_{3}(\xi)$ $=-\operatorname*{sign}(\xi)\xi^{2}+\frac{1}{2}\xi+\frac{1}{8},$
$u_{4}(\xi)$ $=2|\xi|^{3}/3-\xi^{2}/2+1/12,$
$u_{5}(\xi)$ $=-\operatorname*{sign}(\xi)\xi^{4}/3+\xi^{3}/3-\xi/24+1/32\,.$
Thus, we can compute the point evaluations of $f^{a,b}$ and $\widehat{f^{a,b}}$ exactly, both analytically and numerically, and the decay parameters are precisely $a$ and $b$. For the numerical simulations we compute the samples $\widehat{f^{a,b}}(k/p)$ and compare them with the discrete Fourier transform of a sampled version of $f^{a,b}$.
To compute the discrete Fourier transform, we choose $n=2^{l}$, $l=10,\ldots,20$ and then apply the FFT to the samples $f^{a,b}_{h,n}$. The FFT is taken from the Julia package FFTW.jl that provides a binding to the C library FFTW [16].
Since the error estimates of Theorem 2.1 work for every parameter $\alpha<a-1/2$ and $\beta<b-1/2$, we choose $\alpha=a-\frac{1}{2}$ and $\beta=b-\frac{1}{2}$. We compare the results of the numerical experiments with the error rates $n^{-\frac{\alpha\beta}{\alpha+\beta}}$ predicted by Theorem 2.1 for $h=n^{-\frac{\alpha}{\alpha+\beta}}$ that are listed in Table 1.
@@
Our numerical experiments as illustrated in Figure 1 align perfectly with the theoretical bounds in Theorem 2.1 as presented in Table 1 and suggest that they are sharp up to some infinitesimal margin.
Figure 1: Logarithmic plots of $E^{[n]}_{h}(f^{a,b})$ for various $a$ and $b$. Points are computed numerically for $n=2^{l}$, $l=10,\ldots,20$. Reference lines represent the theoretical slopes predicted by Theorem 2.1 and listed in Table 1.
@@
Let us revisit the construction of $f^{a,b}$ to highlight why it is so remarkable that we are able to perform numerical experiments. The B-splines $B_{b}$ and their Fourier transforms $\operatorname*{sinc}^{b}$ can be evaluated accurately. Because B-splines have compact support, the infinite series defining $f^{a,b}$ reduces to a finite sum when evaluated at a finite number of points. Moreover, the careful choice of coefficients in this expansion ensures that the Fourier transform of $f^{a,b}$ admits a closed-form expression and enables the accurate numerical evaluation. This is a rare situation, since most functions do not satisfy these properties, and meaningful numerical experiments depend critically on them.
In particular, the closed-form expressions for $u_{a}(\xi)=\sum_{k\in\mathbb{Z}}c^{a}_{k}e^{-2\pi ik\xi}$ are essential to our numerical results. Without them, the series in (14) would need to be truncated, introducing truncation errors. Summing many terms also risks floating-point inaccuracies, reducing the reliability of computed values of $E^{[n]}_{h}(f^{a,b})$. The availability of a closed-form expression for $u_{a}$ eliminates these sources of numerical error.
Choice of $p$ and $h$
To highlight the importance of a good choice of $p$ and $h$, we consider the function $f^{2,4}$, $f^{2,5}$, $f^{3,5}$, $f^{5,2}$, whose decay in time differs from the decay in frequency. We compare two parameter selections: the optimal choice suggested by Theorem 2.1, versus the standard choice $p=\sqrt{n}$ and $h=1/\sqrt{n}$ [4, 20]. As shown in Figures 2, 3, 4, and 5, the optimal parameters yield an error decay rate consistent with our theoretical predictions and clearly outperform the standard choice.
From the plots one can draw valuable information about the number of samples required to achieve a target accuracy. For instance, by drawing a horizontal line in Figure 2 at the error $E_{h}^{[n]}(f^{2,4})=2^{-15}\approx 3\cdot 10^{-5}$ we find that $n=2^{12}$ samples are required for the optimal choice of $h,n,p$, whereas $2^{17}$ samples are required for the choice $p=h^{-1}=\sqrt{n}$. Even for an accuracy of only five digits, this is an enormous saving!
Figure 2: For $f^{2,4}$, we compare the error curve for the standard choice $p=\sqrt{n}$, $h=1/\sqrt{n}$ with the error curve for the improved choice $p=n^{7/10}$, $h=n^{-3/10}$. The reference lines indicate the theoretical error decay rates: $n^{-3/4}$ for the standard choice and $n^{-21/20}$ for the optimal one.
Figure 3: Comparison of error curves for $f^{2,5}$.
Figure 4: Comparison of error curves for $f^{3,5}$.
Figure 5: Comparison of error curves for $f^{5,2}$.
3.2 Exponential decay in time and polynomial decay in frequency
To demonstrate that Theorem 2.4 is best possible, we consider the exponential functions
$f_{\delta}(x)=e^{-2\pi|x-\delta|},$
for a shift parameter $\delta\in\mathbb{R}$. Its Fourier transform is
$\hat{f_{\delta}}(\xi)=\frac{e^{-2\pi i\delta\xi}}{\pi(1+\xi^{2})}.$
Thus, $f_{\delta}$ decays exponentially in space and polynomially in frequency. This corresponds to the parameters in Theorem 2.4 being $r=2\pi$, $\alpha=1$, and $\sup_{\xi\in\mathbb{R}}|\hat{f_{\delta}}(\xi)|(1+|\xi|)^{b}<\infty$ with $b=2$.
When $p$ is sufficiently large, the error in Theorem 2.4 is dominated by the polynomial decay over a wide range of $n$. For numerical experiments using the FFT with $\delta=0$, we consider the choices
$p=n^{1/2},\quad p=6n^{1/10},\quad\text{and}\quad p=10\,$,
corresponding to $h=n^{-1/2}$, $h=6n^{-9/10}$, and $h=10n^{-1}$, respectively. The expected convergence rates are $h^{3/2}$, which translate into
$n^{-3/4},\quad n^{-27/20},\quad\text{and}\quad n^{-3/2}\,.$
Figure 6 confirms these predictions with excellent agreement between theory and experiment for all shift parameters $\delta$ (the plots look identical for $0\leq\delta\leq 1$).